# Magnitude homology of real hyperplane arrangements

### 04 Apr 2026 v1

@@ 4.5. Magnitude homology on the diagonal
In this subsection, we give a combinatorial formula of the magnitude homology of arrangements on the diagonal $MH_{\ell,\ell}(\mathcal{A})$.
Using Theorem 4.10, we can determine the interior magnitude homology of Boolean arrangements.
@@
The interior magnitude homology of Boolean arrangement $MH_{k,\ell}^{\circ}(Bl_{d})$ concentrates on the diagonal $k=\ell$ and $MH_{\ell,\ell}^{\circ}(Bl_{d})$ is free abelian of rank
@@
We know from Corollary 4.4 and Example 2.14 that $\mathcal{T}(Bl_{d})$ is diagonal and $MH_{\ell,\ell}(Bl_{d})$ is free abelian of rank $\beta_{\ell,\ell}(Bl_{d})=2^{d}\binom{\ell+d-1}{d-1}$. By Theorem 4.10, the interior magnitude homology $MH_{k,\ell}^{\circ}(Bl_{d})$ is a direct summand of $MH_{k,\ell}(Bl_{d})$, and hence concentrates on the diagonal $k=\ell$. Theorem 4.10 also gives
@@
Notice that if $F\in\mathcal{F}(Bl_{d})$ has $\mathrm{codim}F=j$, the localization $(Bl_{d})_{F}$ has essentialization $Bl_{j}$. Then applying Möbius inversion on $\mathcal{F}(\mathcal{A})$ yields the desired formula. ∎
Let us prove the following vanishing result of diagonal interior magnitude homology for non-Boolean arrangements.
@@
Let $\mathcal{A}$ be an arrangement of $\mathrm{rank}\mathcal{A}=r$ and $\mathcal{A}=n$. Suppose $\mathcal{A}$ is not Boolean $r<n$, then the diagonal interior magnitude homology of $\mathcal{A}$ vanishes for all length
$MH_{\ell,\ell}^{\circ}(\mathcal{A})=0,\forall\ell\geq 0.$
@@
Choose an isometric embedding $\phi:\mathcal{T}(\mathcal{A})\hookrightarrow Q_{n}=\{+,-\}^{n}$ of $\mathcal{T}(\mathcal{A})$ into the hypercube $Q_{n}$, which is not surjective since $r<n$. Since $\phi$ is distance preserving, it induces injective chain maps $\phi_{\#}:MC_{*,\ell}(\mathcal{A})\hookrightarrow MC_{*,\ell}(Q_{n})$. Let us take $c\in MC_{\ell,\ell}^{\circ}(\mathcal{A})\cap\ker\partial\subseteq MC_{\ell,\ell}(\mathcal{A})$. Since $\partial c=0$, $\phi_{\#}(c)\in MC_{\ell,\ell}(Q_{n})$ is a cycle. For fixed $\ell$, $MC_{\ell+1,\ell}(Q_{n})=0$, by abuse of notation, we say $\phi_{\#}(c)$ is a homology class in $MH_{\ell,\ell}(Q_{n})$. Recall the isomorphism given in Example 2.14
$\nabla:\bigoplus_{\ell_{1}+\cdots+\ell_{n}=\ell}MH_{\ell_{1},\ell_{1}}(Q_{1})\otimes\cdots\otimes MH_{\ell_{n},\ell_{n}}(Q_{1})\xrightarrow{\cong}MH_{\ell,\ell}(Q_{n}),$
@@
is induced by the Eilenberg–Zilber chain map [16]. We may assume there are basis elements $b^{\lambda}$ of the left hand side, one for each summand, indexed by a composition $\lambda=(\ell_{1},\ldots,\ell_{n})$ of $\ell$, such that $\phi_{\#}(c)=\sum_{\lambda}d_{\lambda}\nabla(b^{\lambda})$ for some scalars $d_{\lambda}$. We must show that $d_{\lambda}=0$.
@@
Since $c\in MC_{\ell,\ell}^{\circ}(\mathcal{A})$, it is a linear combination of walks that cross all hyperplanes in $\mathcal{A}$. Its image $\phi_{\#}(c)$ in $MC_{\ell,\ell}(Q_{n})$ is a linear combination of walks in $Q_{n}$ along which every coordinate gets flipped at least once. This means the basis elements $b^{\lambda}$ involved are indexed by compositions $\lambda=(\ell_{1},\ldots,\ell_{n})$ with all $\ell_{i}\geq 1$. We write such $b^{\lambda}=b_{1}^{\lambda}\otimes\cdots\otimes b_{n}^{\lambda}$, where $b_{i}^{\lambda}\in MH_{\ell_{i},\ell_{i}}(Q_{1})$ is an $\ell_{i}$-step walk in the $i$-th $Q_{1}$. So $b^{\lambda}$ consists of totally $\ell$ steps in these $n$ dimensions with distribution $\lambda$.
@@
By definition of the Eilenberg-Zilber map, $\nabla(b^{\lambda})$ is the signed sum of all possible shuffles of all these $\ell$ steps. It is a linear combination of walks in $Q_{n}$ from the vertex $(s(b_{1}^{\lambda}),\ldots,s(b_{n}^{\lambda}))$ to the vertex $(t(b_{1}^{\lambda}),\ldots,t(b_{n}^{\lambda}))$, where $s(b_{i}^{\lambda})$ and $t(b_{i}^{\lambda})$ represent the starting vertex and terminating vertex in $Q_{1}$ respectively. Since all $\ell_{i}\geq 1$, any vertex in $Q_{n}$ must be visited by at least one walk in $\nabla(b^{\lambda})$. The next key observation is that if $\lambda\neq\mu$, then $\nabla(b^{\lambda})$ and $\nabla(b^{\mu})$ share no common walk due to distribution difference.
Now recall that $\phi:\mathcal{T}(\mathcal{A})\hookrightarrow Q_{n}$ is not surjective, there is a vertex $v_{0}\in Q_{n}$ that is not a vertex of $\mathcal{T}(\mathcal{A})$. We observed that any $\nabla(b^{\lambda})$ contains a walk passing through $v_{0}$ (see Figure 2), but $c\in MC_{\ell,\ell}(\mathcal{A})$ contains no walk passing through $v_{0}$. We must have $d_{\lambda}=0$ in the sum $\sum_{\lambda}d_{\lambda}\nabla(b^{\lambda})=\phi_{\#}(c)$. Therefore we conclude $c=0$, and $MH_{\ell,\ell}^{\circ}(\mathcal{A})=0$.
@@
Figure 2. The tope graph of the $A_{2}$ arrangement embedded as a $C_{6}$ hexagon (blue/bold) inside the Boolean $Q_{3}$ cube. Suppose $b^{\lambda}=(+\xrightarrow{X_{1}}-\xrightarrow{X_{2}}+)\otimes(+\xrightarrow{Y_{1}}-)\otimes(-\xrightarrow{Z_{1}}+)$, where $\lambda=(2,1,1)$. A 4-step shuffle walk $Z_{1}X_{1}X_{2}Y_{1}$ (orange/dotted) in $\nabla(b^{\lambda})$ is shown hitting the forbidden vertex $G^{+}$. Also notice that this walk cannot appear in other $\nabla(b^{\mu})$.
@@
Let $\mathcal{A}$ be an arrangement. The diagonal magnitude homology $MH_{\ell,\ell}(\mathcal{A})$ is free abelian of rank $\beta_{\ell,\ell}(\mathcal{A})$, which is determined by the Boolean flats of the intersection lattice $L(\mathcal{A})$. Precisely, for $\ell\geq 1$,
$\beta_{\ell,\ell}(\mathcal{A})=\sum_{\begin{subarray}{c}X\in L(\mathcal{A})\\ \mathrm{rank}X=\#\mathcal{A}_{X}\end{subarray}}c^{X}\cdot 2^{\mathrm{rank}X}\binom{\ell-1}{\mathrm{rank}X-1}$.
Consequently, $\beta_{\ell,\ell}(\mathcal{A})$ grows as a polynomial in $\ell$ of degree $d_{max}-1$, where $d_{max}=\max\{\mathrm{rank}X\mid X\in L(\mathcal{A}),\mathrm{rank}X=\#\mathcal{A}_{X}\}$.
@@
Since $MC_{\ell+1,\ell}(\mathcal{A})=0$, the diagonal magnitude homology $MH_{\ell,\ell}(\mathcal{A})$ is free abelian. By Theorem 4.10, the global magnitude homology admits a face decomposition. By restricting this decomposition to the diagonal ($k=\ell$) and taking Betti numbers, we obtain
@@
By Theorem 4.14, the interior Betti number $\beta^{\circ}_{\ell,\ell}(\mathcal{A}_{F})$ vanishes unless the localization $\mathcal{A}_{F}$ is Boolean, i.e. $\mathcal{A}_{F}=\mathrm{rank}\mathcal{A}_{F}$. Therefore, the sum collapses to the faces whose support flat $X=s(F)$ satisfies $\mathrm{rank}X=\#\mathcal{A}_{X}$. For these faces, $\mathcal{A}_{F}\cong Bl_{\mathrm{rank}X}$. We can group the non-zero terms of the summation by their support flats $X\in L(\mathcal{A})$. For each valid flat $X$, the number of faces in $\mathcal{A}$ with support $X$ is exactly $c^{X}$. Substituting the known formula for the interior diagonal Betti numbers of the Boolean arrangement (Corollary 4.13), $\beta^{\circ}_{\ell,\ell}(Bl_{d})=2^{d}\binom{\ell-1}{d-1}$ where $d=\mathrm{rank}X$, we obtain
$\beta_{\ell,\ell}(\mathcal{A})=\sum_{\begin{subarray}{c}X\in L(\mathcal{A})\\ \mathrm{rank}X=\#\mathcal{A}_{X}\end{subarray}}c^{X}\cdot 2^{\mathrm{rank}X}\binom{\ell-1}{\operatorname{rank}X-1}$.
For $\ell\geq\mathrm{rank}X$, the binomial $\binom{\ell-1}{\mathrm{rank}X-1}$ is a polynomial in $\ell$ of degree $\mathrm{rank}X-1$. Therefore for $\ell$ sufficiently large, $\beta_{\ell,\ell}(\mathcal{A})$ grows as a polynomial in $\ell$ of degree $d_{max}-1$. ∎
@@ 5. Open problems
Besides the face decomposition conjecture (Conjecture 3.7, see also Corollary 4.11 and Remark 4.12), we propose the following conjectures based on observations of examples in the appendix.
@@
    The coefficients of the power series expression of $\mathrm{Mag}(\mathcal{A},q)$ have alternating signs (equivalently $(-1)^{\ell}\chi_{\ell}(\mathcal{A})\geq 0,\forall\ell\geq 0$).
@@
    $MH_{k,\ell}(\mathcal{A})$ is torsion free.
@@
    If $\mathcal{A}$ is not Boolean, then $\beta_{0,0}(\mathcal{A})=\beta_{r,n}(\mathcal{A})$.
Acknowledgement
The author thanks Weili Guo and Masahiko Yoshinaga for comments on the preliminary draft of this paper.

### 01 May 2026 v3

@@ 5.5. Magnitude homology on the diagonal
In this subsection, we give a combinatorial formula for the diagonal magnitude homology $MH_{\ell,\ell}(\mathcal{A})$ of an arrangement $\mathcal{A}$. First, we determine the interior magnitude homology of Boolean arrangements using Theorem 5.11.
@@
For the Boolean arrangement $Bl_{d}$, the interior magnitude homology $MH_{k,\ell}^{\circ}(Bl_{d})$ is concentrated on the diagonal $k=\ell$, and $MH_{\ell,\ell}^{\circ}(Bl_{d})$ is free abelian of rank
@@
We know from Corollary 5.4 and Example 2.14 that $\mathcal{T}(Bl_{d})$ is diagonal and $MH_{\ell,\ell}(Bl_{d})$ is free abelian of rank $\beta_{\ell,\ell}(Bl_{d})=2^{d}\binom{\ell+d-1}{d-1}$. By Theorem 5.11, the interior magnitude homology $MH_{k,\ell}^{\circ}(Bl_{d})$ is a direct summand of $MH_{k,\ell}(Bl_{d})$, and hence is concentrated on the diagonal $k=\ell$. Theorem 5.11 also gives
@@
If $F\in\mathcal{F}(Bl_{d})$ has $\operatorname{codim}F=j$, then the localization $(Bl_{d})_{F}$ has essentialization $Bl_{j}$. Applying Möbius inversion on $\mathcal{F}(Bl_{d})$ yields the desired formula. ∎
Next we prove the following vanishing result for diagonal interior magnitude homology of non-Boolean arrangements.
@@
Let $\mathcal{A}$ be an arrangement of $\operatorname{rank}\mathcal{A}=r$ and $\mathcal{A}=n$. Suppose $\mathcal{A}$ is not Boolean, so that $r<n$. Then the diagonal interior magnitude homology of $\mathcal{A}$ vanishes in all lengths:
$MH_{\ell,\ell}^{\circ}(\mathcal{A})=0,\quad\forall\ell\geq 0.$
@@
By Definition 5.7, we have
$MH_{\ell,\ell}^{\circ}(\mathcal{A})=\bigoplus_{\begin{subarray}{c}\mathcal{B}\subseteq\mathcal{A}\\ \cap\mathcal{B}=\cap\mathcal{A}\end{subarray}}MH_{\ell,\ell}(\mathcal{A};\mathcal{B}).$
We shall prove that each summand vanishes. In what follows, we suppose that $\mathcal{A}$ is essential without loss of generality, because passing to the essentialization does not alter the tope graph or the intersection lattice.
For $\mathcal{B}\subseteq\mathcal{A}$ such that $\cap\mathcal{B}=\cap\mathcal{A}=\{0\}$, let $m=\#\mathcal{B}$. Then $m\geq r$. Let $\phi:\mathcal{T}(\mathcal{A})\hookrightarrow Q_{\mathcal{A}}=\{+,-\}^{\mathcal{A}}$ be the embedding of $\mathcal{T}(\mathcal{A})$ into the hypercube $Q_{\mathcal{A}}$. Since $\phi$ is distance preserving, it induces an injective chain map $\phi_{\#}:MC_{*,\ell}(\mathcal{A})\hookrightarrow MC_{*,\ell}(Q_{\mathcal{A}})$.
Let $c\in MC_{\ell,\ell}(\mathcal{A};\mathcal{B})$ be a cycle. It maps to a cycle $\phi_{\#}(c)\in MC_{\ell,\ell}(Q_{\mathcal{A}})$. By the endpoint decomposition (2.2), we may suppose without loss of generality that $c$ is a linear combination of $\mathcal{B}$-supported walks with a common starting vertex and a common terminal vertex. This means that the sequence of chambers in each walk of $c$ flips exactly the $m$ coordinates corresponding to $\mathcal{B}$, while the signs on the remaining nmn-m coordinates corresponding to $\mathcal{A}\setminus\mathcal{B}$ remain equal to those of the starting vertex. Geometrically, this means that the entire cycle $\phi_{\#}(c)$ is confined to an $m$-dimensional face of $Q_{\mathcal{A}}$, denoted by $Q_{\mathcal{B}}$ by abuse of notation.
We claim that this subcube $Q_{\mathcal{B}}$ contains a vertex $v_{0}$ that is not a vertex of $\mathcal{T}(\mathcal{A})$. Indeed, in the case $m>r$, the hyperplanes of $\mathcal{B}$ contain linear dependencies. Consequently, the total number of geometrically realizable sign patterns on $\mathcal{B}$ is less than $2^{m}$. However, the subcube $Q_{\mathcal{B}}$ contains all $2^{m}$ sign patterns. Therefore, at least one sign pattern in this subcube is not geometrically realizable, so at least one vertex $v_{0}$ inside the subcube $Q_{\mathcal{B}}$ does not correspond to a chamber in $\mathcal{T}(\mathcal{A})$. In the case $m=r$, $\mathcal{B}$ is a Boolean subarrangement. Its $2^{r}$ chambers form full-dimensional orthants. Because $\mathcal{A}$ is not Boolean, there exists a hyperplane $H\in\mathcal{A}\setminus\mathcal{B}$, and we must have $r\geq 2$. This hyperplane $H$ cannot intersect the interior of every orthant. Thus, there exists a pair of antipodal chambers $C_{v},C_{-v}\in\mathrm{Ch}(\mathcal{B})$ separated by $H$ (that is, $H$ does not cut their interiors), where $v\in\{+,-\}^{\mathcal{B}}$ is the sign vector of $C_{v}$. We may assume that the extension of $v$ to $\tilde{v}\in\{+,-\}^{\mathcal{A}}$ using the signs of the starting vertex of $c$, regarded as a vertex of $Q_{\mathcal{A}}$, is geometrically realizable, i.e. is a vertex of $\mathcal{T}(\mathcal{A})$. Then the extension of v-v to $\widetilde{-v}\in\{+,-\}^{\mathcal{A}}$ is not geometrically realizable, since by definition $\widetilde{-v}_{H}=\widetilde{v}_{H}$, contradicting the geometric condition that $C_{v}$ and $C_{-v}$ are separated by $H$. Thus, we have shown the existence of such a forbidden vertex $v_{0}$ in the subcube $Q_{\mathcal{B}}$.
For a fixed $\ell$, $MC_{\ell+1,\ell}(Q_{\mathcal{A}})=0$; by abuse of notation, we regard $\phi_{\#}(c)$ as a homology class in $MH_{\ell,\ell}(Q_{\mathcal{A}})$. Recall that the isomorphism given in Example 2.14
$\nabla:\bigoplus_{\ell_{1}+\cdots+\ell_{n}=\ell}MH_{\ell_{1},\ell_{1}}(Q_{1})\otimes\cdots\otimes MH_{\ell_{n},\ell_{n}}(Q_{1})\xrightarrow{\cong}MH_{\ell,\ell}(Q_{\mathcal{A}}),$
@@
is induced by the Eilenberg–Zilber chain map [17]. We may assume that there are basis elements $b^{\lambda}$ of the left-hand side, one for each summand, indexed by a composition $\lambda=(\ell_{1},\ldots,\ell_{n})$ of $\ell$, such that $\phi_{\#}(c)=\sum_{\lambda}d_{\lambda}\nabla(b^{\lambda})$ for some scalars $d_{\lambda}$. We must show that $d_{\lambda}=0$.
@@
Since $c$ crosses all hyperplanes in $\mathcal{B}$, its image $\phi_{\#}(c)$ in $MC_{\ell,\ell}(Q_{\mathcal{A}})$ is a linear combination of walks in $Q_{\mathcal{A}}$ along which every $\mathcal{B}$-coordinate is flipped at least once. This means that the basis elements $b^{\lambda}$ involved are indexed by compositions $\lambda=(\ell_{1},\ldots,\ell_{n})$ with all $\ell_{i}\geq 1$ for $i\in\mathcal{B}$. We write such an element as $b^{\lambda}=b_{1}^{\lambda}\otimes\cdots\otimes b_{n}^{\lambda}$, where $b_{i}^{\lambda}\in MH_{\ell_{i},\ell_{i}}(Q_{1})$ is an $\ell_{i}$-step walk in the $i$-th $Q_{1}$. Thus $b^{\lambda}$ consists of $\ell$ steps in total in these $n$ dimensions with distribution $\lambda$.
@@
By definition of the Eilenberg–Zilber map, $\nabla(b^{\lambda})$ is the signed sum of all possible shuffles of all these $\ell$ steps. It is a linear combination of walks in $Q_{\mathcal{A}}$ from the vertex $(s(b_{1}^{\lambda}),\ldots,s(b_{n}^{\lambda}))$ to the vertex $(t(b_{1}^{\lambda}),\ldots,t(b_{n}^{\lambda}))$, where $s(b_{i}^{\lambda})$ and $t(b_{i}^{\lambda})$ represent the starting vertex and terminal vertex in $Q_{1}$, respectively. Since all $\ell_{i}\geq 1$ for $i\in\mathcal{B}$, any vertex in the subcube $Q_{\mathcal{B}}$ must be visited by at least one walk in $\nabla(b^{\lambda})$. The next key observation is that if $\lambda\neq\mu$, then $\nabla(b^{\lambda})$ and $\nabla(b^{\mu})$ share no common walk because their step distributions differ.
@@
Figure 3. The tope graph of the $A_{2}$ arrangement embedded as a $C_{6}$ hexagon (blue/bold) inside the Boolean $Q_{3}$ cube. Suppose $b^{\lambda}=(+\xrightarrow{X_{1}}-\xrightarrow{X_{2}}+)\otimes(+\xrightarrow{Y_{1}}-)\otimes(-\xrightarrow{Z_{1}}+)$, where $\lambda=(2,1,1)$. A 4-step shuffle walk $Z_{1}X_{1}X_{2}Y_{1}$ (orange/dotted) in $\nabla(b^{\lambda})$ is shown passing through the forbidden vertex $G^{+}$. Note also that this walk cannot appear in any other $\nabla(b^{\mu})$.
Now we apply the above observations to the forbidden vertex $v_{0}$. Each $\nabla(b^{\lambda})$ contains a walk passing through $v_{0}$ (see Figure 3), but $c\in MC_{\ell,\ell}(\mathcal{A})$ contains no walk passing through $v_{0}$. We must have $d_{\lambda}=0$ in the sum $\sum_{\lambda}d_{\lambda}\nabla(b^{\lambda})=\phi_{\#}(c)$. Therefore $c=0$ and $MH_{\ell,\ell}(\mathcal{A};\mathcal{B})=0$. ∎
@@
Let $\mathcal{A}$ be an arrangement. The diagonal magnitude homology $MH_{\ell,\ell}(\mathcal{A})$ is free abelian of rank $\beta_{\ell,\ell}(\mathcal{A})$, which is determined by the Boolean flats of the intersection lattice $L(\mathcal{A})$. More precisely, for $\ell\geq 1$,
$\beta_{\ell,\ell}(\mathcal{A})=\sum_{\begin{subarray}{c}X\in L(\mathcal{A})\\ \operatorname{rank}X=\#\mathcal{A}_{X}\end{subarray}}c^{X}\cdot 2^{\operatorname{rank}X}\binom{\ell-1}{\operatorname{rank}X-1}$.
Consequently, $\beta_{\ell,\ell}(\mathcal{A})$ grows as a polynomial in $\ell$ of degree $d_{\max}-1$, where
$d_{\max}=\max\{\operatorname{rank}X\mid X\in L(\mathcal{A}),\ \operatorname{rank}X=\#\mathcal{A}_{X}\}.$
@@
Since $MC_{\ell+1,\ell}(\mathcal{A})=0$, the diagonal magnitude homology $MH_{\ell,\ell}(\mathcal{A})$ is free abelian. By Theorem 5.11, the global magnitude homology admits a face decomposition. Restricting this decomposition to the diagonal ($k=\ell$) and taking Betti numbers, we obtain
@@
By Theorem 5.16, the interior Betti number $\beta^{\circ}_{\ell,\ell}(\mathcal{A}_{F})$ vanishes unless the localization $\mathcal{A}_{F}$ is Boolean, i.e. $\mathcal{A}_{F}=\operatorname{rank}\mathcal{A}_{F}$. Therefore, the sum collapses to the faces whose support flat $X=s(F)$ satisfies $\operatorname{rank}X=\#\mathcal{A}_{X}$. For these faces, $\mathcal{A}_{F}\cong Bl_{\operatorname{rank}X}$. We can group the nonzero terms of the summation by their support flats $X\in L(\mathcal{A})$. For each valid flat $X$, the number of faces in $\mathcal{A}$ with support $X$ is exactly $c^{X}$. Substituting the known formula for the interior diagonal Betti numbers of the Boolean arrangement (Corollary 5.15), namely $\beta^{\circ}_{\ell,\ell}(Bl_{d})=2^{d}\binom{\ell-1}{d-1}$ where $d=\operatorname{rank}X$, we obtain
$\beta_{\ell,\ell}(\mathcal{A})=\sum_{\begin{subarray}{c}X\in L(\mathcal{A})\\ \operatorname{rank}X=\#\mathcal{A}_{X}\end{subarray}}c^{X}\cdot 2^{\operatorname{rank}X}\binom{\ell-1}{\operatorname{rank}X-1}$.
For $\ell\geq\operatorname{rank}X$, the binomial $\binom{\ell-1}{\operatorname{rank}X-1}$ is a polynomial in $\ell$ of degree $\operatorname{rank}X-1$. Therefore, for $\ell$ sufficiently large, $\beta_{\ell,\ell}(\mathcal{A})$ grows as a polynomial in $\ell$ of degree $d_{\max}-1$. ∎
@@ 6. Open problems
Based on observations from examples in the appendix, we propose the following conjectures.
@@
    $(-1)^{\ell}\chi_{\ell}(\mathcal{A})\geq 0$ for sufficiently large $\ell$.
@@
    $MH_{k,\ell}(\mathcal{A})$ is torsion-free.
@@
    Suppose that $r\geq 2$ and that $\mathcal{A}$ is indecomposable, that is, $\mathcal{A}$ cannot be expressed as a direct sum $\mathcal{A}_{1}\oplus\mathcal{A}_{2}$ of nonzero arrangements. Then $\beta_{0,0}=\beta_{r,n}$.
Acknowledgements
The authors thank Weili Guo, Yoh Kitajima, Takuya Saito, and Masahiko Yoshinaga for their comments on a preliminary version of this paper.